# RoHOI: Robustness Benchmark for Human-Object Interaction Detection

### v1

@@ 4 Method
Score Guided Progressive Feature Learning. To fully exploit the potential of the proposed SAM-guided semantic masks for maximizing HOI cue extraction, we introduce a novel score-guided progressive feature learning framework that incorporates a frequency memory bank across all severity levels.
Initially, training begins with the lowest severity level, while a memory bank is designed to track the utilization frequency of different severity levels. The perturbation level is then progressively increased to a more challenging level based on the relative change in the evaluation metric between consecutive epochs. Notably, the severity level is only upgraded when the selected level differs from $\omega_{1}$. Let $\Delta Q(t)$ represent the relative change in the evaluation score on the validation set at epoch $t$:
@@
5 Experimental Results
5.1 Implementation Details
[…]
5.2 Analysis of the Benchmark
@@
Figure 2: Comparison of SAMPL with strong two-stage and one-stage baselines trained on V-COCO under RoHOI benchmarks. The left bar chart groups performance by corruption categories, while the right radar chart compares models across five severity levels for three corruptions. Corruptions are denoted as Abbreviation, i.e., $L\gamma$, where $\gamma$ represents the severity level. Details are shown in Sec. 3.1.
@@
Table 1 presents the performance comparison between SAMPL and existing one-stage and two-stage HOI detection methods on HICO-DET [62] and V-COCO [13] datasets. SAMPL surpasses RLIPv2 [70], improving MRI by $3.78$ and CRI by $12.5\%$ on V-COCO, demonstrating superior robustness against visual degradations. On HICO-DET, SAMPL achieves a $3.6\%$ higher CRI while maintaining comparable MRI, ensuring strong generalization without compromising clean-set accuracy. These results are further supported by Fig. 2, where SAMPL consistently outperforms two-stage and one-stage baselines across all $4$ corruption categories in the RoHOI benchmark, achieving the highest MRI in each group and demonstrating improvements over the current SOTA model at every severity level (detailed in the supplementary). While extreme noise levels still pose challenges, SAMPL significantly improves robustness compared to prior methods. These results highlight SAMPL’s effectiveness in enhancing HOI detection under corrupted conditions through our proposed semantic-aware masking and progressive learning. Through a comprehensive analysis of representative HOI models, we derive the following key findings.
@@
Observation 1: Large-scale vision-language pretraining significantly enhances the robustness of transformer-based models. According to the experiments conducted both on the clean test set and our perturbed test set on the two datasets, shown in Table 1, we observe that transformer-based architectures exhibit varying levels of robustness depending on their pretraining strategies. RLIPv2 [70], which leverages large-scale multimodal pretraining, achieves substantially higher robustness scores in contrast to GEN-VLKT [36], SOV-STG [6] and other transformer-based methods. This suggests that large-scale vision-language pretraining provides a more structured relational feature space, mitigating feature drift and localization sensitivity in HOI detection tasks. Furthermore, two-stage models that utilize pretrained detector backbone, e.g., UPT [74], show only moderate robustness improvements, highlighting the importance of diverse relational supervision beyond object-centric datasets.
@@
Figure 3: Qualitative comparison of HOI detection under corruption between SAMPL (ours) and RLIPv2 [70]. The first row shows uncorrupted images, while the second and third rows present SAMPL and RLIPv2 results on corrupted images. Missed predictions are in dashed red, false interactions in red, and correct ones in blue.
@@
Observation 2: Object localization strategies dictate robustness in two-stage architectures. Two-stage methods show varying robustness across different object localization strategies. Region-proposal-based models [72] achieve high clean-set accuracy but degrade under corruption due to error propagation in localization. In contrast, query-based methods [74] adapt object representations dynamically, enhancing robustness on HICO-DET. However, this advantage diminishes on V-COCO, where UPT with a larger backbone exhibits lower CRI than CDN [72] with a smaller backbone, likely due to V-COCO’s limited interaction diversity. This suggests that query-based detection benefits from diverse datasets but struggles with constrained interaction distributions.
@@
RLIPv2-ParSeDA$+$SAMSM 68.27 70.24 48.20 0.27
RLIPv2-ParSeDA$+$AGNP 68.33 70.51 44.74 0.24
RLIPv2-ParSeDA$+$RPP 67.69 69.64 47.59 0.25
@@
Observation 3: Interaction representation entanglement amplifies robustness degradation under perturbations. Models with tightly coupled interaction representations achieve higher clean-set accuracy but are more prone to corruption due to feature distortion propagation. MUREN [27], with multi-level relational context learning, suffers the largest MRI and CRI drop, showing that excessive entanglement weakens robustness. Similarly, GEN-VLKT-L’s global self-attention leads to greater degradation than CDN-L’s decoupled interaction classification, while QAHOI’s [5] adaptive query-based approach improves robustness by dynamically refining predictions. This underscores a key trade-off: stronger feature coupling boosts accuracy but reduces resilience to corruptions.
@@
5.3 Analysis of the Ablation Study
To investigate the impact of perturbation-based training on model robustness, we conduct a series of ablation experiments on the V-COCO dataset. The results, presented in Tables 2 and 3, analyze the effects of different perturbation types and perturbation level settings on model performance.
@@ 5.3 Analysis of the Ablation Study
Figure 4: Performance degradation of RLIPv2 [70] on HICO-DET [62] and V-COCO [13] under RoHOI. The left radar chart shows the relative mAP (Full) drop on HICO-DET, while the right shows the decline in $\mathrm{AP}^{\#2}_{\text{role}}$ on V-COCO. Clean dataset performance serves as a reference, with shaded areas indicating corruption impact, with abbreviations mentioned in Sec. 3.1.
Effect of Perturbation Levels: Table 3 analyzes the impact of masking severity on robustness. SAMSM occludes $40\%$$80\%$ of each instance’s bounding box, with the $40\%/50\%/60\%$ setting achieving the best trade-off, maximizing MRI and CRI while maintaining competitive AP scores. Moderate occlusions enhance feature learning via contextual cues, while excessive masking ($>70\%$) disrupts spatial structures, reducing discriminative capacity. These findings emphasize the need for controlled occlusions that challenge models without excessive information loss.
@@
5.4 Analysis of the Qualitative Results
Fig. 3 compares SAMPL with RLIPv2 [70] under various corruptions, demonstrating SAMPL’s superior robustness in the first four cases while highlighting shared failure patterns in the last two. Under Screen Crack corruption (Column 1), SAMPL accurately detects occluded background persons and their “ride” interactions, whereas RLIPv2 misses smaller subjects, indicating SAMPL’s stronger resilience to fragmented textures. In Pixelation noise (Column 2), our method correctly recognizes “work on computer”, while RLIPv2 misclassifies it as “cut”, suggesting greater robustness against structured perturbations. Under Salt-and-Pepper Noise (Column 3), RLIPv2 confuses “sit” with “ride”, likely due to the presence of an adjacent animal, whereas SAMPL maintains stable interaction recognition, reducing reliance on object-centric biases. In Moire Pattern (Column 4) corruption, RLIPv2 incorrectly detects “talk on phone”, possibly due to moire-induced distortions interfering with high-frequency texture recognition. However, under high-severity Glass Blur (Column $5$) and Moire Pattern corruption (Column 6), both models misclassify actions as “surf” or “talk on phone”, potentially because V-COCO’s limited action categories encourage overfitting to common interactions under extreme degradation rather than a traditional long-tail distribution issue. These results highlight SAMPL’s advantages in structured occlusions and texture-based corruptions. However, both models remain vulnerable under severe degradations, where dataset priors dominate predictions.
6 Conclusion
@@ Appendix B Standard HOI Evaluation Metrics
This section provides a detailed explanation of the evaluation metrics used for HICO-DET [64] and V-COCO [13] in our main paper. These include the standard mAP metrics for HICO-DET and AProle variations for V-COCO.
[…]
@@
B.2 $\mathrm{AP}^{\#1}_{\text{role}}$ and $\mathrm{AP}^{\#2}_{\text{role}}$
@@
CDN-L [72] 15.38 10.12 48.73 7.56 41.67 60.92 28.78 56.89 24.24 28.18 64.67 41.06 49.75 67.07 30.67 44.30 37.33 50.79 22.39 22.45
GEN-VLKT-L [36] 14.64 9.90 48.27 7.74 40.50 60.63 28.37 56.31 20.40 28.52 64.94 40.48 50.48 67.48 32.42 44.21 38.27 51.15 21.11 23.24
MUREN [27] 15.71 12.10 51.08 6.77 41.65 62.59 26.08 56.32 20.46 23.89 66.65 38.55 51.78 69.12 25.63 44.36 38.79 52.34 15.97 22.22
RLIPv2-ParSeDA [70] 25.31 17.84 57.82 12.99 50.91 66.72 50.58 64.68 40.11 38.34 68.39 55.99 54.61 69.82 45.93 59.14 48.14 53.01 26.12 29.73
SAMPL (Ours) 26.63 19.03 58.56 13.87 52.46 68.06 52.66 65.81 42.24 46.69 69.22 57.31 62.04 70.77 47.89 60.72 48.99 53.21 29.90 30.49
@@
Motion Blur (MB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Defocus Blur (DB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Gaussian Blur (GauB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Glass Blur (GB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 5: Visual examples of Optical System (OS)-Induced artifacts across severity levels.
@@ Appendix C Corruption Types in the Dataset
C.1 Optical System-Induced (OS) Artifacts
Motion Blur (MB): Blur caused by the relative motion between the camera and the scene during exposure. This artifact is included as it commonly occurs in dynamic scenes or when handheld cameras are used. Severity levels are generated by increasing the kernel size of a linear motion blur filter. (Details shown in Table 6, row $1$.)
Defocus Blur (DB): Blur resulting from the lens being improperly focused on the subject. This noise reflects the limitations of camera focus systems, particularly in low-light or high-speed scenarios. Severity levels are controlled by varying the radius of a circular point spread function (PSF). (Details shown in Fig. 6, row $2$.)
Gaussian Blur (GauB): Blur due to scattering or inherent optical limitations of the lens. This artifact simulates the diffusion of light caused by imperfect optical surfaces or aperture settings. Severity levels are adjusted by increasing the standard deviation of the Gaussian kernel. (Details shown in Fig. 6, row $3$.)
Glass Blur (GB): Blur introduced when imaging through imperfect or contaminated transparent materials (e.g., glass or plastic). This type is included to represent common real-world challenges such as photographing through windows or protective covers. Severity levels are generated by applying iterative small displacements and local blurring. (Details shown in Fig. 6, row $4$.)
@@ C.1 Optical System-Induced (OS) Artifacts
Salt-and-Pepper Noise (S&P) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
JPEG Artifact (JPEG) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Speckle Noise (SN) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Gaussian Noise (GauN) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Shot Noise (ShN) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Packet Loss (PL) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 6: Visual examples of Sensor, Compression and Transmission (SCT) artifacts across severity levels.
C.2 Sensor, Compression and Transmission (SCT) Artifacts
@@
Salt-and-Pepper Noise (S&P): Discrete black-and-white pixel noise caused by sensor malfunction or transmission errors. This represents rare but impactful noise patterns in hardware-limited imaging systems. Severity levels are generated by increasing the probability of random pixel corruption. (Details shown in Fig. 6, row $1$.)
@@
JPEG Artifacts (JPEG): Blocky or blurry artifacts resulting from lossy image compression, which are used to evaluate model robustness to low-quality image formats frequently encountered in data storage and transmission. Severity levels are controlled by progressively reducing the JPEG quality factor. (Details shown in Fig. 6, row $2$.)
@@
Speckle Noise (SN): Speckle noise resembles the grainy texture seen on old printed photos or low-quality digital images taken in dim lighting. At lower severity levels, it appears as faint specks of dust on a camera lens, barely noticeable unless viewed closely. As severity increases, the noise becomes more pronounced, resembling the rough texture of sandpaper or the graininess of a poorly printed newspaper. At the highest severity levels, the effect is akin to heavy rain blurring a car windshield, significantly degrading image clarity. (Details shown in Fig. 6, row $3$.)
@@
Gaussian Noise (GauN): Additive noise caused by sensor imperfections such as thermal noise or quantization errors. This is a ubiquitous noise model, reflecting baseline sensor limitations in most imaging devices. Severity levels are determined by progressively increasing the noise standard deviation. (Details shown in Fig. 6, row $4$.)
@@
Shot Noise (ShN): Random variations in pixel intensity caused by photon-counting statistics during exposure. This noise type represents the fundamental physical limits of image sensors in low-light conditions. Severity levels are controlled by adjusting the scale of the Poisson-distributed noise. (Details shown in Fig. 6, row $5$.)
@@
Packet Loss (PL): Missing or distorted image regions caused by data corruption during transmission. This noise simulates real-world challenges in wireless or network-based imaging systems. Severity levels are generated by increasing the number and size of corrupted or duplicated regions. (Details shown in Fig. 6, row $6$.)
C.3 Environmentally Induced (EI) Artifacts
@@
Exposure (EXP) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Rainbow Effect (RE) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Occlusion (OCC) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Vignette Effect (VE) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 7: Visual examples of Environmentally Induced (EI) artifacts across severity levels.
Exposure Artifacts (EXP): Loss of image details caused by underexposure or overexposure due to extreme lighting conditions. Higher severity levels increase the intensity of overexposure (bright regions saturate) or underexposure (dark regions lose details), leading to stronger visual degradation. This type is included to represent real-world challenges in scenes with poor or uncontrolled lighting. (Details shown in Fig. 7, row $1$.)
Rainbow Effect (RE): Typically caused by the interaction between the projection system’s optical components (e.g., color wheels in DLP projectors) and the camera sensor. As severity increases, the color distortion becomes more pronounced, with broader and more intense rainbow bands appearing across the image, significantly altering local color distribution. This noise enables the evaluation of model robustness against chromatic distortions, particularly in scenes involving projected displays and reflective surfaces. (Details shown in Fig. 7, row $2$.)
Occlusion (OCC): Partial obstruction of the field of view by external objects. Increasing severity levels introduce randomly placed black rectangular occlusions of varying sizes, progressively removing more contextual information from the image. This is a critical inclusion as occlusions are prevalent in crowded or dynamic environments. (Details shown in Fig. 7, row $3$.)
Vignette Effect (VE): Gradual darkening at the edges of an image caused by lens design or aperture settings. With higher severity, the vignette effect has darker and wider peripheral regions, increasingly suppressing details in the outer parts of the image. This type represents a common optical limitation that affects peripheral details in wide-field imaging. (Details shown in Fig. 7, row $4$.)
C.4 Geometric and Scene (G&S) Distortions
@@
Moire Pattern (MP) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Screen Crack (SC) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Elastic Transform (ET) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Perspective Distortion (PD) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pixelation (PIX) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Zoom Blur (ZB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 8: Visual examples of Geometric and Scene (G&S) Distortions across severity levels.

### v2

@@ Introduction
[…]
Score Guided Progressive Feature Learning. To effectively utilize SAM-guided masks, we design a novel score-guided progressive learning framework that incorporates a frequency memory bank across severity levels.
Initially, training begins with the lowest severity level, while a memory bank is designed to track the utilization frequency of different severity levels. The perturbation level is then progressively increased based on the relative change in the evaluation metric between consecutive epochs. Notably, the severity level is only upgraded when the selected level differs from $\omega_{1}$. Let $\Delta Q(t)$ represent the relative change in the evaluation score on the validation set at epoch $t$:
@@
Experimental Results
Implementation Details. Our SAMPL method is built upon RLIPv2 (Yuan et al. 2023) with a Swin-Tiny (Liu et al. 2021) backbone and has $214M$ trainable parameters. The backbone is pretrained on Visual Genome (Krishna et al. 2017), MS-COCO (Lin et al. 2014) and Object365 (Shao et al. 2019), and fine-tuned on the HICO-DET or V-COCO datasets. We use the AdamW optimizer (Loshchilov and Hutter 2019) with an initial learning rate of 1ee-4 (backbone and text encoder at 1ee-5), linear warmup followed by cosine decay, weight decay of 1ee-4, and batch size $2$ per GPU with gradient accumulation over $2$ steps. Models are trained for $80$ epochs using cross-entropy loss for object classification, focal loss for verb classification, and L1 + GIoU loss for bounding, with KL divergence for latent space regularization. The encoder and decoder consist of $6$ and $3$ layers, respectively. For the dynamic severity adjustment mechanism, we set the initial threshold as $\tau_{\text{init}}=0.15$ and $\epsilon=10^{-6}$.
@@
Figure 2: Comparison of SAMPL with strong two-stage and one-stage baselines trained on V-COCO under RoHOI benchmarks. The left bar chart groups performance by corruption categories, while the right radar chart compares models across five severity levels for three corruptions. Corruptions are denoted as Abbreviation, i.e., $L\gamma$, where $\gamma$ represents the severity level. Details are shown in Sec. Benchmark.
@@
Analysis of the Benchmark. Table 1 presents the performance comparison between SAMPL and existing one-stage and two-stage HOI detection methods on HICO-DET (Wang et al. 2020) and V-COCO (Gupta and Malik 2015) datasets. SAMPL surpasses RLIPv2 (Yuan et al. 2023), improving MRI by $3.78$ and CRI by $12.5\%$ on V-COCO, demonstrating superior robustness against visual degradations. On HICO-DET, SAMPL achieves a $3.6\%$ higher CRI while maintaining comparable MRI, ensuring strong generalization without compromising clean-set accuracy. These results are further supported by Fig. 2, where SAMPL consistently outperforms two-stage and one-stage baselines across all $4$ corruption categories in the RoHOI benchmark, achieving the highest MRI in each group and demonstrating improvements over the current SOTA model at every severity level (detailed in the supplementary). While extreme noise levels still pose challenges, SAMPL significantly improves robustness compared to prior methods by progressively training the model with semantic-aware occlusions, enabling it to learn context-driven features that are resilient to diverse and severe corruptions. These results highlight SAMPL’s effectiveness in enhancing HOI detection under corrupted conditions through our proposed semantic-aware masking and progressive learning. Through a comprehensive analysis of representative HOI models, we derive the following key findings.
@@
Observation 1: Large-scale vision-language pretraining significantly enhances the robustness of transformer-based models. Experiments on the clean and corrupted test set (Table 1) reveal that transformer-based architectures exhibit robustness variations dependent on their pretraining strategies. RLIPv2, which leverages large-scale multimodal pretraining, achieves substantially higher robustness scores in contrast to GEN-VLKT, SOV-STG and other transformer-based methods. This suggests that large-scale vision-language pretraining provides a more structured relational feature space, mitigating feature drift and localization sensitivity in HOI detection tasks. Furthermore, two-stage models that utilize pretrained detector backbone, e.g., UPT, show only moderate robustness improvements, highlighting the importance of diverse relational supervision beyond object-centric datasets.
@@
Figure 3: Qualitative comparison of HOI detection under corruption between SAMPL (ours) and RLIPv2 (Yuan et al. 2023). The first row shows uncorrupted images, while the second and third rows present SAMPL and RLIPv2 results on corrupted images. Missed predictions are in dashed red, false interactions in red, and correct ones in blue.
@@
Observation 2: Object localization strategies dictate robustness in two-stage architectures. Two-stage methods show varying robustness across different object localization strategies. Region-proposal-based models (Zhang et al. 2021) achieve high clean-set accuracy but degrade under corruption due to error propagation in localization. In contrast, query-based methods (Zhang, Campbell, and Gould 2022) adapt object representations dynamically, enhancing robustness on HICO-DET. However, on V-COCO, UPT with a larger backbone surprisingly shows lower CRI than CDN with a smaller backbone, likely attributed to V-COCO’s limited interaction diversity. This suggests that query-based detection benefits from diverse datasets but struggles with constrained interaction distributions.
@@
[…]
@@
Observation 3: Interaction representation entanglement amplifies robustness degradation under perturbations. Models with tightly coupled interaction representations achieve higher clean-set accuracy but are more prone to corruption due to feature distortion propagation. MUREN, with multi-level relational context learning, suffers the largest MRI and CRI drop, showing that excessive entanglement weakens robustness. Similarly, GEN-VLKT-L’s global self-attention leads to greater degradation than CDN-L’s decoupled interaction classification, while QAHOI’s adaptive query-based approach improves robustness by dynamically refining predictions. This underscores a key trade-off: stronger feature coupling boosts accuracy but reduces resilience to corruptions.
@@
Analysis of the Ablation Study. To investigate the impact of perturbation-based training on model robustness, we conduct a series of ablation experiments on the V-COCO dataset. The results, presented in Tables 2 and 3, analyze the effects of different perturbation types and perturbation level settings on model performance.
@@
Figure 4: Performance degradation of RLIPv2 (Yuan et al. 2023) on HICO-DET and V-COCO under RoHOI. The left radar chart shows the relative mAP (Full) drop on HICO-DET, while the right shows the decline in $\mathrm{AP}^{\#2}_{\text{role}}$ on V-COCO. Clean dataset performance serves as a reference, with shaded areas indicating corruption impact.
Effect of Perturbation Levels: Table 3 analyzes the impact of masking severity on robustness. SAMSM occludes $40\%$$80\%$ of each instance’s bounding box, with the $40\%/50\%/60\%$ setting achieving the best trade-off, maximizing MRI and CRI while maintaining competitive AP scores. Moderate occlusions enhance feature learning via contextual cues, while excessive masking ($70\%$) disrupts spatial structures, reducing discriminative capacity. These findings emphasize the need for controlled occlusions that challenge models without excessive information loss.
@@
Analysis of the Qualitative Results. Fig. 3 compares SAMPL with RLIPv2 under various corruptions, demonstrating SAMPL’s superior robustness in the first four cases while highlighting shared failure patterns in the last two. Under Screen Crack corruption (Col. $1$), SAMPL accurately detects occluded background persons and their “ride” interactions, whereas RLIPv2 misses smaller subjects, indicating SAMPL’s stronger resilience to fragmented textures, as SAMPL uses semantic-aware masking that selectively occludes HOI-relevant regions, training the model to recognize interactions from partial and degraded visual cues. In Pixelation noise (Col. $2$), our method correctly recognizes “work on computer”, while RLIPv2 misclassifies it as “cut”, suggesting greater robustness against structured perturbations. Under Salt-and-Pepper Noise (Col. $3$), RLIPv2 confuses “sit” with “ride”, likely due to the presence of an adjacent animal, whereas SAMPL maintains stable interaction recognition, reducing reliance on object-centric biases. In Moire Pattern (Col. $4$) corruption, RLIPv2 incorrectly detects “talk on phone”, possibly due to moire-induced distortions interfering with high-frequency texture recognition. However, under high-severity Glass Blur (Col. $5$) and Moire Pattern corruption (Col. $6$), both models misclassify actions as “surf” or “talk on phone”, potentially because V-COCO’s limited action categories encourage overfitting to common interactions under extreme degradation rather than a traditional long-tail distribution issue. These results highlight SAMPL’s advantages in structured occlusions and texture-based corruptions. However, both models remain vulnerable under severe degradations, where dataset priors dominate predictions.
Conclusion
@@ Appendix B Standard HOI Evaluation Metrics
This section provides a detailed explanation of the evaluation metrics used for HICO-DET (Xu et al. 2019) and V-COCO (Gupta and Malik 2015) in our main paper. These include the standard mAP metrics for HICO-DET and AProle variations for V-COCO.
$\bm{\mathrm{mAP}}\,\textbf{(Full, Rare, Non-Rare)}$
@@
$\bm{\mathrm{AP}^{\#1}_{\text{role}}}$ and $\bm{\mathrm{AP}^{\#2}_{\text{role}}}$
@@
CDN-L (Zhang et al. 2021) 15.38 10.12 48.73 7.56 41.67 60.92 28.78 56.89 24.24 28.18 64.67 41.06 49.75 67.07 30.67 44.30 37.33 50.79 22.39 22.45
GEN-VLKT-L (Liao et al. 2022) 14.64 9.90 48.27 7.74 40.50 60.63 28.37 56.31 20.40 28.52 64.94 40.48 50.48 67.48 32.42 44.21 38.27 51.15 21.11 23.24
MUREN (Kim, Jung, and Cho 2023) 15.71 12.10 51.08 6.77 41.65 62.59 26.08 56.32 20.46 23.89 66.65 38.55 51.78 69.12 25.63 44.36 38.79 52.34 15.97 22.22
RLIPv2-ParSeDA (Yuan et al. 2023) 25.31 17.84 57.82 12.99 50.91 66.72 50.58 64.68 40.11 38.34 68.39 55.99 54.61 69.82 45.93 59.14 48.14 53.01 26.12 29.73
[…]
@@
Motion Blur (MB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Defocus Blur (DB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Gaussian Blur (GauB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Glass Blur (GB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Visual examples of Optical System (OS)-Induced artifacts across severity levels.
@@ Appendix C Corruption Types in the Dataset
Optical System-Induced (OS) Artifacts
Motion Blur (MB): Blur caused by the relative motion between the camera and the scene during exposure. This artifact is included as it commonly occurs in dynamic scenes or when handheld cameras are used. Severity levels are generated by increasing the kernel size of a linear motion blur filter. (Details shown in Figure 5, row $1$.)
Defocus Blur (DB): Blur resulting from the lens being improperly focused on the subject. This noise reflects the limitations of camera focus systems, particularly in low-light or high-speed scenarios. Severity levels are controlled by varying the radius of a circular point spread function (PSF). (Details shown in Figure 5, row $2$.)
Gaussian Blur (GauB): Blur due to scattering or inherent optical limitations of the lens. This artifact simulates the diffusion of light caused by imperfect optical surfaces or aperture settings. Severity levels are adjusted by increasing the standard deviation of the Gaussian kernel. (Details shown in Figure 5, row $3$.)
Glass Blur (GB): Blur introduced when imaging through imperfect or contaminated transparent materials (e.g., glass or plastic). This type is included to represent common real-world challenges such as photographing through windows or protective covers. Severity levels are generated by applying iterative small displacements and local blurring. (Details shown in Figure 5, row $4$.)
@@
Salt-and-Pepper Noise (S&P) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
JPEG Artifact (JPEG) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Speckle Noise (SN) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Gaussian Noise (GauN) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Shot Noise (ShN) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Packet Loss (PL) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 6: Visual examples of Sensor, Compression and Transmission (SCT) artifacts across severity levels.
Sensor, Compression and Transmission (SCT) Artifacts
@@
Salt-and-Pepper Noise (S&P): Discrete black-and-white pixel noise caused by sensor malfunction or transmission errors. This represents rare but impactful noise patterns in hardware-limited imaging systems. Severity levels are generated by increasing the probability of random pixel corruption. (Details shown in Figure 6, row $1$.)
@@
JPEG Artifacts (JPEG): Blocky or blurry artifacts resulting from lossy image compression, which are used to evaluate model robustness to low-quality image formats frequently encountered in data storage and transmission. Severity levels are controlled by progressively reducing the JPEG quality factor. (Details shown in Figure 6, row $2$.)
@@
Speckle Noise (SN): Speckle noise resembles the grainy texture seen on old printed photos or low-quality digital images taken in dim lighting. At lower severity levels, it appears as faint specks of dust on a camera lens, barely noticeable unless viewed closely. As severity increases, the noise becomes more pronounced, resembling the rough texture of sandpaper or the graininess of a poorly printed newspaper. At the highest severity levels, the effect is akin to heavy rain blurring a car windshield, significantly degrading image clarity. (Details shown in Figure 6, row $3$.)
@@
Gaussian Noise (GauN): Additive noise caused by sensor imperfections such as thermal noise or quantization errors. This is a ubiquitous noise model, reflecting baseline sensor limitations in most imaging devices. Severity levels are determined by progressively increasing the noise standard deviation. (Details shown in Figure 6, row $4$.)
@@
Shot Noise (ShN): Random variations in pixel intensity caused by photon-counting statistics during exposure. This noise type represents the fundamental physical limits of image sensors in low-light conditions. Severity levels are controlled by adjusting the scale of the Poisson-distributed noise. (Details shown in Figure 6, row $5$.)
@@
Packet Loss (PL): Missing or distorted image regions caused by data corruption during transmission. This noise simulates real-world challenges in wireless or network-based imaging systems. Severity levels are generated by increasing the number and size of corrupted or duplicated regions. (Details shown in Figure 6, row $6$.)
Environmentally Induced (EI) Artifacts
@@
Exposure (EXP) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Rainbow Effect (RE) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Occlusion (OCC) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Vignette Effect (VE) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Visual examples of Environmentally Induced (EI) artifacts across severity levels.
Exposure Artifacts (EXP): Loss of image details caused by underexposure or overexposure due to extreme lighting conditions. Higher severity levels increase the intensity of overexposure (bright regions saturate) or underexposure (dark regions lose details), leading to stronger visual degradation. This type is included to represent real-world challenges in scenes with poor or uncontrolled lighting. (Details shown in Figure 7, row $1$.)
Rainbow Effect (RE): Typically caused by the interaction between the projection system’s optical components (e.g., color wheels in DLP projectors) and the camera sensor. As severity increases, the color distortion becomes more pronounced, with broader and more intense rainbow bands appearing across the image, significantly altering local color distribution. This noise enables the evaluation of model robustness against chromatic distortions, particularly in scenes involving projected displays and reflective surfaces. (Details shown in Figure 7, row $2$.)
Occlusion (OCC): Partial obstruction of the field of view by external objects. Increasing severity levels introduce randomly placed black rectangular occlusions of varying sizes, progressively removing more contextual information from the image. This is a critical inclusion as occlusions are prevalent in crowded or dynamic environments. (Details shown in Figure 7, row $3$.)
Vignette Effect (VE): Gradual darkening at the edges of an image caused by lens design or aperture settings. With higher severity, the vignette effect has darker and wider peripheral regions, increasingly suppressing details in the outer parts of the image. This type represents a common optical limitation that affects peripheral details in wide-field imaging. (Details shown in Figure 7, row $4$.)
Geometric and Scene (G&S) Distortions
@@
Moire Pattern (MP) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Screen Crack (SC) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Elastic Transform (ET) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Perspective Distortion (PD) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Pixelation (PIX) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Zoom Blur (ZB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 8: Visual examples of Geometric and Scene (G&S) Distortions across severity levels.